OpenStudy (anonymous):

A rower travels upstream at 6 km per hour and back to the starting place at 10 km per hour. The total journey takes 48 minutes. How far upstream did the rower go?

Directrix (directrix):

Suggested solution courtesy of @czar ----------- Use the equation d = rt where d = distance, r = rate or speed, and t = time. Solve for d upstream. d + d = r(up & down) * t (total). d / 6 + d / 10 = 48/60 hours. Multiply all terms by common denominator of 60 to cancel out denominators. 10d + 6d = 48 16d = 48. d = ?

Directrix (directrix):

@hannah.grace97